OpenStudy (anonymous):

ax+bx-1=8 solve for x

OpenStudy (anonymous):

factorize

OpenStudy (akashdeepdeb):

ax+bx-1=8 ax + bx = 8+1 x(a+b) = 9 x = 9/(a+b) Understood? :)

OpenStudy (anonymous):

thanks @AkashdeepDeb

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

x=9